OpenStudy (anonymous):

I need to evaluate the expression. But i'm not sure on the steps to solve it. Can you help? 12 divided by (7-3)^2+2

OpenStudy (katrinakaif):

First (7-3) = 4 Square 4 = 16 and add 2 to get 18 on the bottom 12/18 and simplify

OpenStudy (anonymous):

Solve (7-3) first = 4. 4^2 = 16 + 2 = 18. Denominator is 18. 12/18.. Simplify.

OpenStudy (anonymous):

Oh...Ok. Thank you for explaining it so well. :)

OpenStudy (anonymous):

Oh...Ok. Thank you for explaining it so well. :)